OpenStudy (linn99123):

neeed help fast will fan and give medal Which set of integers is not a Pythagorean triple? A. 9, 40, 41 B. 11, 60, 61 C. 15, 18, 21 D. 20, 21, 29

OpenStudy (whpalmer4):

a Pythagorean triple \(a,b,c\) has the property that \(a^2+b^2=c^2\) The best known triple is probably \(3,4,5\)

OpenStudy (whpalmer4):

A trick you could use to find which ones might work or not without having to square all the numbers: the last digit of the square is the last digit of the number you are squaring, squared. So 41^2 = some number ending in 1. If we square 9, that ends in 1. If we square 40, that ends in 0. 0+1=1 so the first one might work. If we ended up with something like 1+3, then we would know that was not a triple.

OpenStudy (whpalmer4):

so if \(15,17,19\) was on the list: last digit of 15^2 = 5*5 = 25 so 5 last digit of 17^2 = 7*7 = 49 so 9 last digit of 19^2 = 9*9 = 81 so 1 5+9 = 14, which doesn't have 1 as its final digit, so we would know that \(15,17,19\) is not a triple.

OpenStudy (anonymous):

eliminate A, B, and D 15^2=225 18^2=324 225+324=549 21^2=441 549=441 ( N/A) it doesnt work So C is your answer

OpenStudy (anonymous):

C